# S2ST-Omni: An Efficient and Scalable Multilingual Speech-to-Speech Translation Framework via Seamlessly Speech-Text Alignment and Streaming Speech Decoder

### v1

@@ Abstract
Multilingual speech-to-speech translation (S2ST) aims to directly convert spoken utterances from multiple source languages into natural and intelligible speech in a target language. Despite recent progress, significant challenges remain: 1) achieving high-quality and low-latency S2ST remains a critical hurdle; 2) existing S2ST approaches heavily rely on large-scale parallel speech corpora, which are extremely difficult to collect. To address these issues, we propose S2ST-Omni, an efficient and scalable framework for multilingual speech-to-speech translation. To be specific, we decompose the S2ST task into speech-to-text translation (S2TT) and text-to-speech synthesis (TTS), unifying them within a single end-to-end speech-language model. To achieve high-quality S2TT while reducing dependence on parallel corpora, we leverage large-scale pretrained models—Whisper for audio understanding and Qwen 3.0 for text understanding. A lightweight speech adapter is further introduced to seamlessly align speech and text representations, enabling effective utilization of the pretrained multimodal knowledge. To ensure both translation quality and real-time performance, we adopt a pretrained streaming speech decoder in the TTS stage, generating target speech in an autoregressive manner. Extensive experiments on the CVSS benchmark demonstrate that S2ST-Omni outperforms state-of-the-art S2ST baselines while maintaining comparable latency, highlighting its effectiveness and practical potential for real-world deployment.
@@ 4.1.2 Implementation Details
In all experiments, we employ the AdamW optimizer to train the proposed S2ST-Omni model, with an initial learning rate set to $1\times 10^{-5}$. The batch sizes for the speech-to-text translation (S2TT) and text-to-speech (TTS) stages are 32 and 128, respectively. All models are trained for 20 epochs using the PyTorch framework on four NVIDIA A100 GPUs. The system prompt for the Qwen 3.0 large language model (LLM) is defined as: “You are an expert in speech-to-speech translation, responsible for accurately translating multilingual input speech into corresponding English text.”
@@
4https://dl.fbaipublicfiles.com/fairseq/wav2vec/wav2vec_vox_960h_pl.pt5https://github.com/mjpost/sacrebleu
@@ 4.2 Results and Analysis
To validate the effectiveness of the proposed method, we compare the translation performance of S2ST-Omni against several state-of-the-art S2ST systems. A comprehensive summary of the results is presented in Table 1.
[…]
@@
In summary, the experimental results provide strong evidence for the effectiveness and practicality of S2ST-Omni in multilingual speech-to-speech translation.

### v2

# S2ST-Omni: An Efficient and Scalable Speech-to-Speech Translation Framework via Seamlessly Speech-Text Alignment and Streaming Speech Decoder

@@ Abstract
Multilingual speech-to-speech translation (S2ST) aims to directly convert spoken utterances from multiple source languages into fluent and intelligible speech in a target language. Despite recent progress, several critical challenges persist: 1) achieving high-quality and low-latency S2ST remains a significant obstacle; 2) most existing S2ST methods rely heavily on large-scale parallel speech corpora, which are difficult and resource-intensive to obtain. To tackle these challenges, we introduce S2ST-Omni, a novel, efficient, and scalable framework tailored for multilingual speech-to-speech translation. To enable high-quality S2TT while mitigating reliance on large-scale parallel speech corpora, we leverage powerful pretrained models: Whisper for robust audio understanding and Qwen 3.0 for advanced text comprehension. A lightweight speech adapter is introduced to bridge the modality gap between speech and text representations, facilitating effective utilization of pretrained multimodal knowledge. To ensure both translation accuracy and real-time responsiveness, we adopt a streaming speech decoder in the TTS stage, which generates the target speech in an autoregressive manner. Extensive experiments conducted on the CVSS benchmark demonstrate that S2ST-Omni consistently surpasses several state-of-the-art S2ST baselines in translation quality, highlighting its effectiveness and superiority.
@@ 4.1.2 Implementation Details
In all experiments, we employ the AdamW optimizer to train the proposed S2ST-Omni model, with an initial learning rate set to $1\times 10^{-5}$. The batch sizes for the speech-to-text translation (S2TT) and text-to-speech (TTS) stages are 32 and 128, respectively. All models are trained for 100,000 steps using the PyTorch framework on four NVIDIA A100 GPUs. The system prompt for the used Qwen 3.0 4B LLM is defined as: “You are an expert in speech-to-speech translation, responsible for accurately translating multilingual input speech into corresponding English text.”
@@
4https://dl.fbaipublicfiles.com/fairseq/wav2vec/wav2vec_vox_960h_pl.pt5https://github.com/mjpost/sacrebleu6https://facebookresearch.github.io/stopes/docs/eval/blaser
@@ 4.2 Results and Analysis
To validate the effectiveness of the proposed approach, we compare the translation performance of S2ST-Omni against six state-of-the-art S2ST systems. A comprehensive summary of the results is provided in Table 1.
As shown in the table, S2ST-Omni consistently and significantly outperforms all baselines on the CVSS benchmark across the evaluated language pairs, with gains exceeding 38 BLEU points in certain cases. These results highlight the strong generalization capability of the proposed framework across linguistically diverse settings. Specifically, in the French $\rightarrow$ English direction, S2ST-Omni achieves an ASR-BLEU score of 67.45, markedly surpassing the strongest baseline (StreamSpeech, 28.45). This substantial improvement underscores the model’s robustness in handling phonetic variability and prosodic features characteristic of French. For German $\rightarrow$ English, the model attains a remarkable 71.20 ASR-BLEU score—the highest among all language pairs—demonstrating its effectiveness in managing syntactic complexity and compound structures inherent to German. In the Spanish $\rightarrow$ English setting, S2ST-Omni achieves a strong 68.28 ASR-BLEU score, again outperforming all baselines by a wide margin. This result reflects the model’s ability to capture the rich morphological patterns of Spanish and generate fluent, accurate translations.
The outstanding performance of S2ST-Omni can be attributed to several key architectural innovations:
@@
Overall, the substantial and consistent improvements across all language pairs validate the robustness and efficacy of the proposed S2ST-Omni framework, establishing it as a new state-of-the-art in direct speech-to-speech translation with promising potential for real-world multilingual applications.